OpenStudy (anonymous):

Find the area of a sector defined by a central angle of 120 degrees and a radius of 6 cm. a) 3*pi square cm b) 6*pi square cm c) 12*pi square cm d) 36*pi square cm I got answer C. Is that right? Thank you to anyone who can help!

OpenStudy (anonymous):

\[\frac{ Your Angle}{ 360 } \pi r^2\]In our case r=6 and you r angle =120

OpenStudy (anonymous):

So it's 120/360 times 36*pi?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

so it's C. :)

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

thank you :)